The amount of an investment of P dollars for t years at simple rate r is given by A = p + prt
A) Rewrite this formula by factoring ourt the greatest common factor on the right hand side.
B) Find A if $8300 is invested for 3 years at a simple rate of 15%.

Found 2 solutions by Earlsdon, stanbon:
Answer by Earlsdon(6294)

1) Factor a p from
A = p + prt
A = p(1+rt)
2) Find A if p = $8,300, r = 0.15, and t = 3
A = $8,300(1+(0.15)(3))
A = $8,300(1.45)
A = $12,035

Answer by stanbon(75887)

The amount of an investment of P dollars for t years at simple rate r is given by A = p + prt
A) Rewrite this formula by factoring out the greatest common factor on the right hand side.
A = p(1+rt)
----------------------------
B) Find A if $8300 is invested for 3 years at a simple rate of 15%.
A = 8300(1+0.15*3)
A = 8300(1.45)
A = 12035
----------------------------
Cheers,
Stan H.